# Generative modeling of conditional probability distributions on the level-sets of collective variables

### v2

@@ 4.3 Alanine Dipeptide
In this study, we applied our generative approach to alanine dipeptide, which is a small benchmark molecular system consisting of 22 atoms. Our goal is to train the ODE model (10) for generative modeling of configurations of the system’s 10 non-Hydrogen atoms conditioned on its two backbone dihedral angles $\phi$ and $\psi$.
@@
Trajectory data of the alanine dipeptide system were generated by performing molecular dynamics simulations for $t=1500\,\textrm{ns}$ using the GROMACS package [47], where we adopted the same simulation parameters as in the previous work [24] and the states were stored every $10$ ps, resulting in a trajectory dataset of $1.5\times 10^{5}$ states. Only Cartesian coordinates of the 10 non-Hydrogen atoms were then extracted using MDAnalysis [12, 37] and aligned to a reference configuration via the Kabsch algorithm [20], thereby eliminating global translational and rotational motions. The aligned coordinates were then standardized by subtracting the empirical mean and dividing by the empirical standard deviation before being used as training data. Figure 12(a) shows the empirical density of the trajectory data projected onto the space of the two dihedral angles.
@@
In addition to the unbiased MD data, we also generated an independent trajectory by performing an ABF simulation for $t=500\,\textrm{ns}$ along the backbone dihedral angles $\phi$ and $\psi$. The ABF method applies a time-dependent biasing force to enhance the sampling of rarely visited regions of conformational space. The simulation was performed using the ColVars module [11] and the GROMACS package. The states were stored every $10$ steps and the resulting trajectory dataset consists of $5\times 10^{4}$ states. Figure 12(b) shows the empirical density of the trajectory data projected onto the space of the two dihedral angles. All ABF-biased coordinates were preprocessed in the same manner as the unbiased data—centered, aligned to a reference structure, flattened, and then standardized.
For the modeling of the vector field $v$ in (10), we implemented a fully connected feedforward neural network that takes as input the atomic coordinates $x\in\mathbb{R}^{30}$, the temporal parameter $t\in[0,1]$, and the CV value $z\in\mathbb{R}^{2}$, which correspond to the standardized values of $(\phi,\psi)$. The feedforward network consists of five hidden layers with 512 hidden units and Tanh activations, while the output layer returns the vector field in $\mathbb{R}^{30}$.
The training of the vector field $v$ was performed by minimizing the objective (13) for 1000 epochs, using the Adam optimizer with a learning rate of $10^{-3}$, a batch size of 512, and default PyTorch parameters (e.g., for momentum coefficients, numerical stability parameter, and weight decay). After training, new samples in $\mathbb{R}^{30}$ can be generated conditioned on the value of $z$ (i.e. the target value of $(\phi,\psi)$) by integrating the ODE (10) from random initial positions sampled from the standard Gaussian distribution. The generated samples are then denormalized and reshaped to obtain three-dimensional coordinates of the $10$ non-Hydrogen atoms.
Results.
To assess the quality of the generated conformations, we chose the target dihedral angles $(\phi,\psi)\in\{(-75^{\circ},150^{\circ}),(52.49^{\circ},34.95^{\circ}),(57.5^{\circ},-126.78^{\circ})\}$, which correspond respectively to the densely populated region, a secondary metastable region, and an isolated point in a low-population region in the dihedral angle space (see Figure 12(a)). For each of these values $(\phi,\psi)$, we generated 1000 samples and identified the (reference) configuration in the trajectory whose dihedral angles are closest to $(\phi,\psi)$. Root mean square deviation (RMSD) values of the generated samples from the reference were computed after Kabsch alignment. Histograms of RMSD values were constructed with fixed bin widths ($0.02$ Å) to visualize the distribution of deviations in different regions represented by these values of dihedral angles. These distributions are shown for the model trained using the unbiased dataset in Figure 13 and for the model trained using the biased dataset in Figure 14, respectively.
RMSD distributions between nearby real structures are shown for the unbiased trajectory dataset in Figures 15(a) and 15(b), serving as a baseline for intrinsic structural variability. In the left panel (Figure 15(a)), all pairs of configurations whose dihedral angles $(\phi_{1},\psi_{1})$ and $(\phi_{2},\psi_{2})$ satisfy $(\phi_{1}-\phi_{2})^{2}+(\psi_{1}-\psi_{2})^{2}<1^{\circ}$ were identified using a nearest neighbors search in the dihedral angle space. For each pair, the RMSD was computed after Kabsch alignment to remove global translations and rotations. The right panel (Figure 15(b)) extends this analysis by binning the $\phi$ and $\psi$ angles of configurations into a $50\times 50$ grid and computing the mean RMSD of all conformations within each bin (empty bins were ignored). For the ABF-biased dataset, the same evaluation workflow was applied, and the results were nearly identical to those for the unbiased dataset. Comparing the histograms in Figure 13 and Figure 15, we can conclude that the mean deviations predicted by the trained model are of the same order as the intrinsic structural variability observed in the unbiased dataset.
To assess the angular consistency of the generated configurations across the dihedral space, we computed heatmaps of mean deviations (see (24)) per $(\phi,\psi)$-bin for the generative models trained using both the unbiased and the ABF-biased datasets, as shown in Figure 16. For this analysis, the $\phi$ and $\psi$ angles in the range $180^{\circ},180^{\circ}]$ were binned into intervals of equal length $20^{\circ}$, and the centers of each bin were used as target CV values. For each target CV value, $N=200$ configurations were generated using the trained ODE models. The resulting mean deviations are shown in Figure 16(a) for the model trained using the unbiased trajectory data, where small deviations are concentrated in the high-probability regions of the CV space, and in Figure 16(b) for the model trained using the ABF-biased trajectory data, which exhibit uniformly low deviations across the full CV space.
Similarly to the previous example, the projection method introduced in Section 3.3 can be applied as a post-processing step to modify the generated configurations so that they satisfy the constraints on the target dihedral angles. We refer to the previous work [34, Appendices B.2 and B.4] for related numerical study.
@@ Results.
Figure 12: Empirical densities of the dihedral angles of the trajectory data. (a) Trajectory from unbiased simulation. (b) Trajectory from ABF simulation.
Figure 13: Results of the generative model trained with the unbiased dataset. Distributions of RMSD values (in Å) for generated samples at three representative CV locations: (a) $(\phi,\psi)=(-75^{\circ},150^{\circ})$ in a well-populated region (b) $(\phi,\psi)=(52.49^{\circ},34.95^{\circ})$ in the secondary metastable basin, and (c) $(\phi,\psi)=(57.5^{\circ},-126.78^{\circ})$ corresponding to an isolated point in the low-population region.
Figure 14: Results of the generative model trained with the ABF-biased dataset. Distributions of RMSD values (in Å) for generated samples at the same representative CV locations as in Figure 13: (a) $(\phi,\psi)=(-75^{\circ},150^{\circ})$, (b) $(\phi,\psi)=(52.49^{\circ},34.95^{\circ})$, and (c) $(\phi,\psi)=(57.5^{\circ},-126.78^{\circ})$.
Figure 15: Analysis of intrinsic structural variability in the unbiased real dataset. (a) Distribution of RMSD values (in Å) between pairs of nearby configurations satisfying $(\phi_{1}-\phi_{2})^{2}+(\psi_{1}-\psi_{2})^{2}<1$. (b) Histogram of mean RMSD values computed for each $50\times 50$ $(\phi,\psi)$-bin in the dihedral angle space, capturing spatial variations in local conformational variability.
@@
Figure 16: Heat maps of the mean deviation between the target CV value $(\phi,\psi)$ and the dihedral angles of generated configurations. (a) Unbiased molecular dynamics data. (b) ABF-biased data. Each pixel represents the mean circular distance in degrees, quantifying the accuracy with which the conditional generative model reproduces the desired angular constraints. Lower deviation indicates higher fidelity of the generated ensemble to the target CVs.
@@ 5 Conclusion
In this work, we have introduced a general framework for the generative modeling of conditional distributions on level-sets of CVs using flow-matching neural ODEs. By conditioning the learned flows on CV values, the model can generate samples across multiple level-sets without requiring explicit geometric information. To further enhance accuracy on level-sets in low-probability regions, we proposed to leverage biased trajectory data from enhanced sampling techniques such as ABF. Validation on concrete datasets demonstrates that our approach accurately reproduces conditional distributions and successfully exploits biased data to improve learning in scarcely sampled regions.
Beyond these results, this study opens several promising directions for future work. First, to better assess the effectiveness of the approach, it is meaningful to study the accuracy of the generative model by comparing certain distances between the generated distributions and the conditional distributions on level-sets. Second, a natural extension would be to combine the approach proposed in this work with methods that propagate the effective (reduced) dynamics in the CV space [21, 50]. Such a combination would lead to generative modeling approaches that can generate new trajectories of stochastic processes while taking into account the essential dynamics.
@@ 6 Acknowledgements
Fatima-Zahrae Akhyar acknowledges the Zuse institute Berlin for hosting her internship during which this research was conducted. Wei Zhang has been funded through DFG’s “Eigene Stelle” program (project 524086759) and has been supported by the DFG-grant CRC 1114 “Scaling Cascades in Complex Systems” Project Number 235221301. Gabriel Stoltz was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project EMC2, grant agreement No 810367), and was supported by Hi! PARIS and ANR/France 2030 program (ANR-23-IACL-0005). Christof Schütte has been supported through DFG-grant CRC 1114 (Project A04 “Efficient calculation of slow and stationary scales in molecular dynamics”, Project A05 “Probing scales in equilibrated systems by optimal nonequilibrium forcing”, Project B03 “Multilevel coarse graining of multiscale problems”, and Project C03 “Multiscale modelling and simulation for spatiotemporal master equation”).

### v3

@@ 4.3 Alanine Dipeptide
We applied our generative approach to alanine dipeptide, which is a small benchmark molecular system consisting of 22 atoms. We studied two different generative modeling tasks. In the first task, our goal is to train the ODE model (10) to generate configurations of the system’s 10 non-Hydrogen atoms conditioned on its two backbone dihedral angles $\phi$ and $\psi$. In the second task, our goal is to train the ODE model (10) to generate configurations of the system’s 10 non-Hydrogen atoms conditioned on the dihedral angle $\phi$ alone.
@@
Trajectory data of the alanine dipeptide system in water was generated by performing molecular dynamics simulations for $t=1500\,\textrm{ns}$ using the GROMACS package [47], where we adopted the same simulation parameters as in the previous work [24] and the states were recorded every $10$ ps, resulting in a trajectory dataset of $1.5\times 10^{5}$ states. Cartesian coordinates of the 10 non-Hydrogen atoms were then extracted using MDAnalysis [12, 37] and aligned to a reference configuration via the Kabsch algorithm [20], thereby eliminating global translational and rotational motions. The aligned coordinates were then standardized by subtracting the empirical mean and dividing by the empirical standard deviation before being used as training data. Figure 12(a) shows the empirical density of the trajectory data projected onto the space of the two dihedral angles.
@@
In addition to the unbiased trajectory, we also generated two biased trajectory datasets of the system by performing ABF simulations using the ColVars module [11] and the GROMACS package. Specifically, for the first task (i.e. the generation conditioned on both dihedral angles $\phi$ and $\psi$), we performed an ABF simulation for $t=500\,\textrm{ns}$, where the CV was defined as the two dihedral angles $\phi$ and $\psi$. The states were stored every $10$ ps and the resulting trajectory dataset consists of $5\times 10^{4}$ states. For the second task (i.e. the generation conditioned on the dihedral angle $\phi$ alone), we ran an ABF simulation for $t=100\,\textrm{ns}$, where the CV was defined as the backbone dihedral angle $\phi$ alone. The states were stored every $1$ ps, resulting in a trajectory dataset of $10^{5}$ states. Figures 12(b) and 12(c) show the empirical densities of the two trajectory datasets projected onto the space of the two dihedral angles, respectively. It can be observed that different values of the corresponding CVs are more uniformly sampled under the ABF simulations. Figures 13(a) and 13(b) display the free energy profiles associated to the two corresponding CVs (reconstructed using the ColVars module), but this information on free energies is not used in the study below. Finally, coordinates of states in both datasets were preprocessed in the same manner as the unbiased data—centered, aligned to a reference structure, flattened, and then standardized.
@@
Figure 12: Empirical densities of the dihedral angles of the trajectory data. (a) Trajectory from the unbiased simulation. (b) Trajectory from the ABF simulation, where the CV is defined as the dihedral angles $(\phi,\psi)$. (c) Trajectory from the ABF simulation, where the CV is defined as the dihedral angle $\phi$.
Figure 13: (a) Free energy profile associated to the two dihedral angles $(\phi,\psi)$. (b) Free energy profile associated to the dihedral angle $\phi$.
For the modeling of the vector field $v$ in (10), we implemented a fully connected feedforward neural network that takes as input the atomic coordinates $x\in\mathbb{R}^{30}$, the temporal parameter $t\in[0,1]$, and a value $z$ that corresponds to the conditioning (either angles or their trigonometric functions). Specifically, in the first task, the value $z$ is defined either as the standardized values of $(\phi,\psi)$ in $\mathbb{R}^{2}$, or as $z=(\cos\phi,\sin\phi,\cos\psi,\sin\psi)\in\mathbb{R}^{4}$. In the second task, $z$ is defined either as the standardized value of $\phi$ in $\mathbb{R}$, or as $z=(\cos\phi,\sin\phi)\in\mathbb{R}^{2}$. The advantage of using trigonometric functions is that the vector field is guaranteed to be periodic as a function of angles. In all cases, the feedforward network consists of five hidden layers with 512 hidden units and Tanh activations, while the output layer returns the vector field in $\mathbb{R}^{30}$.
The training of the vector field $v$ was performed by minimizing the objective (13) for 1000 epochs, using the Adam optimizer with a learning rate of $10^{-3}$, a batch size of 512, and default PyTorch parameters (e.g. momentum coefficients, numerical stability parameter, and weight decay). For each of the two tasks, the training was carried out both on the unbiased dataset and on the ABF-biased dataset. After training, new samples in $\mathbb{R}^{30}$ can be generated conditioned on the target CV value (i.e. the target value of $(\phi,\psi)$ in the first task, or the target value of $\phi$ in the second task) by integrating the ODE (10) from random initial positions sampled from the standard Gaussian distribution. The generated samples are then denormalized and reshaped to obtain three-dimensional coordinates of the $10$ non-Hydrogen atoms.
Results for the first task.
To assess the quality of the generated configurations, we chose the target dihedral angles $(\phi,\psi)\in\{(-70^{\circ},150^{\circ}),(52.5^{\circ},35^{\circ})\}$, which correspond to a state in the densely populated region and a state in the secondary metastable region, respectively (see Figure 12(a)). For each of these values $(\phi,\psi)$, a configuration in the unbiased trajectory whose dihedral angles are closest to the target value $(\phi,\psi)$ was identified as a reference, and 2000 samples were generated conditioned on the target angles for each of the trained models (i.e. models trained on the unbiased dataset and on the ABF-biased dataset, with conditioning $z$ being either angles or trigonometric functions). The root mean square deviation (RMSD) values of the generated configurations from the reference (after Kabsch alignment) were computed. Figures 14(a)14(b) show the estimated densities of the RMSDs. Moreover, we selected the configurations in the dataset whose dihedral values belong to the same cell as the target dihedral values $(\phi,\psi)$, when binned into the uniform grid of $180^{\circ},180^{\circ}]\times[-180^{\circ},180^{\circ}]$ with width $10^{\circ}$. The RMSD values of the selected configurations from the reference were computed and the estimated densities are also shown in Figures 14(a)14(b) for comparison. It can be observed that the density profiles of RMSDs computed for the generated configurations are similar to the density profiles computed for the configurations selected from the trajectory datasets, indicating that the trained models are able to produce consistent configurations conditioned on the target dihedral angles. The fact that the density profiles are unimodal is also consistent with the fact that the configuration of alanine dipeptide is largely determined by its two dihedral angles.
To assess the angular consistency of the generated configurations across the dihedral angle space, we computed heat maps of mean deviations (see (24)) per $(\phi,\psi)$-bin for the trained models, as shown in Figure 15. For this analysis, the $\phi$ and $\psi$ angles in the domain $180^{\circ},180^{\circ}]\times[-180^{\circ},180^{\circ}]$ were binned using a uniform grid of equal width $20^{\circ}$ (in both directions), and the centers of grid cells were used as target CV values. For each target CV value, $N=200$ configurations were generated using the trained ODE models. Figure 15(a) and Figure 15(b) show the resulting mean deviations for the models trained on the unbiased trajectory data, where the standardized dihedral angles and the trigonometric functions $(\cos\phi,\sin\phi,\cos\psi,\sin\psi)\in\mathbb{R}^{4}$ were passed to the neural network, respectively. For both models, small deviations are concentrated in the high-probability regions of the CV space. Comparing these two figures, it is clear that using trigonometric functions is advantageous, as it leads to lower mean deviations in the region close to the right boundary $\phi=180^{\circ}$ in the CV space, since the data points close to the left boundary $\phi=-180^{\circ}$ can contribute to the learning in this region thanks to the periodicity of the vector field (see Figure 12(a)). Similarly, Figures 15(c) and 15(d) show the mean deviations for the models trained on the ABF-biased trajectory data, where the standardized dihedral angles and their trigonometric functions were used as inputs of the neural network, respectively. Comparing to the models trained on the unbiased trajectory data, lower deviations are observed across the CV space for the models trained on the ABF-biased trajectory data. Since in this case the CV space is well covered by the ABF-biased trajectory data, only slight improvement is achieved when trigonometric functions are used (Figure 15(d)).
Figure 14: Density estimations of RMSD values (in Å) for two target dihedral angles: (a) $(\phi,\psi)=(-70^{\circ},150^{\circ})$ and (b) $(\phi,\psi)=(52.5^{\circ},35^{\circ})$ in the first task. In both figures, the line “unbiased, ref” (resp. “biased, ref”) corresponds to the density of RMSDs computed for the configurations in the unbiased (resp. ABF-biased) trajectory dataset whose dihedral angles are close to the target dihedral values (see the text for details). The line with label “unbiased” (resp. “biased”) corresponds to the model trained on the unbiased (resp. ABF-biased) dataset, where the standardized angles are directly passed to the neural network. The line with label “unbiased, tri” (resp. “biased, tri”) corresponds to the model trained on the unbiased (resp. ABF-biased) dataset, where the values $(\cos\phi,\sin\phi,\cos\psi,\sin\psi)$ are passed to the neural network.
@@
Figure 15: Heat maps of the mean deviation (i.e. the mean circular distance in degrees) between the target CV value $(\phi,\psi)$ and the dihedral angles of generated configurations in the first task. (a) Model trained on the unbiased trajectory data, where the standardized dihedral angles are used as input of the neural network. (b) Model trained on the unbiased trajectory data, where the trigonometric functions $(\cos\phi,\sin\phi,\cos\psi,\sin\psi)\in\mathbb{R}^{4}$ are used as input of the neural network. (c) Model trained on the ABF-biased trajectory data, where the standardized dihedral angles are used as input of the neural network. (d) Model trained on the ABF-biased trajectory data, where the trigonometric functions $(\cos\phi,\sin\phi,\cos\psi,\sin\psi)\in\mathbb{R}^{4}$ are used as input of the neural network. Lower deviation indicates higher fidelity of the generated ensemble to the target CVs.
Results for the second task.
Regarding the generative modeling task conditioned on the angle $\phi$ alone, we tested the models trained on the unbiased trajectory data (Figure 12(a)) and on the ABF-biased data (Figure 12(c)), and for each case the vector field $v$ takes either the standardized angle $\phi$ or the trigonometric functions $(\cos\phi,\sin\phi)\in\mathbb{R}^{2}$ as inputs for the conditioning.
The interval $180^{\circ},180^{\circ}]$ was divided into subintervals of equal length $20^{\circ}$ and the centers of the subintervals were chosen as target CV values of $\phi$. For each of these target values and for each of the trained models, $200$ samples were generated, based on which the mean deviations (see (24)) were computed. The results are shown in Figure 16(a). Similar to the results in the first task, when the unbiased trajectory data is used in training, low deviations are achieved for values of $\phi$ that are frequently taken by the configurations in the dataset, whereas the deviations are high for values of $\phi$ that are rarely taken by configurations in the dataset. In this case, the use of trigonometric functions $(\cos\phi,\sin\phi)\in\mathbb{R}^{2}$ helps to reduce the deviations for target $\phi$ values that are close to $180^{\circ}$. When the ABF-biased trajectory data is used in training, the model accuracy is greatly improved, as the deviations are low across different target values $\phi$. In this case, the advantage of using trigonometric functions is less obvious since the entire space of $\phi$ is well covered by the data (Figure 12(c)).
To further assess the quality of the models, we studied the distribution of $\psi$ conditioned on the target values $\phi=-70^{\circ}$ and $\phi=52.5^{\circ}$. For each of these target values and for each of the trained models, 1000 samples were generated, based on which the empirical density of the angle $\psi$ was estimated. For comparison, the density of $\psi$ conditioned on the target value $\phi$ was also estimated using configurations in the unbiased trajectory dataset and configurations in the ABF-biased trajectory dataset (by selecting configurations whose angle $\phi$ belongs to the same cell as the target value of $\phi$ when binned using a uniform grid with width $20^{\circ}$). The results are shown in Figures 16(b) and 16(c). As can be observed there, for each of the trained models, the distribution of the angle $\psi$ of the generated samples matches well with the conditional distributions estimated using configurations in the trajectory datasets.
Finally, as in the previous example, the projection method introduced in Section 3.3 can be applied as a post-processing step to modify the generated configurations so that they satisfy the constraints on the target dihedral angles exactly. We refer to the previous work [34, Appendices B.2 and B.4] for a related numerical study.
Figure 16: Results for the second task. (a) Mean deviation (in degrees) as a function of the target value $\phi$. (2) Estimated densities of $\psi$ conditioned on $\phi=-70^{\circ}$. (3) Estimated densities of $\psi$ conditioned on $\phi=52.5^{\circ}$. In each figure, the line with label “unbiased” (resp. “biased”) corresponds to the model trained on the unbiased (resp. ABF-biased) trajectory dataset, where the model takes the standardized value of $\phi$ as an input. The line with label “unbiased, tri” (resp. “biased, tri”) corresponds to the model trained on the unbiased (resp. ABF-biased) trajectory dataset, where the model takes the trigonometric functions $(\cos\phi,\sin\phi)\in\mathbb{R}^{2}$ as input. In (b) and (c), the line with labels “unbiased, ref” (resp. “biased, ref”) corresponds to the density of $\psi$ estimated using configurations in the unbiased (resp. ABF-biased) trajectory dataset.
@@ 5 Conclusion
In this work, we have introduced a general framework for the generative modeling of conditional distributions on level-sets of CVs using flow-matching neural ODEs. By conditioning the learned flows on CV values, the model can generate samples across different level-sets without requiring explicit geometric information. To enhance the generation accuracy on level-sets in low-probability regions, we proposed to leverage biased trajectory data from enhanced sampling techniques such as ABF. Validation on concrete datasets demonstrates that our approach accurately reproduces conditional distributions and successfully exploits biased data to improve learning in scarcely sampled regions.
This study opens several promising directions for future work. First, it is interesting to apply the method to more complex molecular systems and study how the method performs (in terms of the amount of required data and the model quality) when the size of the molecular system increases, especially in the case where the target conditional distributions are multimodal. For such applications, it could be the case that more advanced architectures (e.g. using featurization or transformers) would improve the generative capabilities. Second, a natural extension would be to combine the approach proposed in this work with methods for identifying good CV maps and methods for propagating the effective (reduced) dynamics in the CV space [21, 50]. Such a combination would lead to generative modeling approaches that can generate new trajectories of high-dimensional complex stochastic dynamics while taking the essential dynamics into account.
@@ 6 Acknowledgements
Fatima-Zahrae Akhyar acknowledges Zuse institute Berlin for hosting her internship during which this research was conducted. Wei Zhang has been funded through DFG’s “Eigene Stelle” program (project 524086759) and has been supported by the Young Investigator Program of the DFG-grant CRC 1114 “Scaling Cascades in Complex Systems” Project Number 235221301. Gabriel Stoltz was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project EMC2, grant agreement No 810367), and was supported by Hi! PARIS and ANR/France 2030 program (ANR-23-IACL-0005). Christof Schütte has been supported through DFG-grant CRC 1114 (Project A04 “Efficient calculation of slow and stationary scales in molecular dynamics”, Project A05 “Probing scales in equilibrated systems by optimal nonequilibrium forcing”, Project B03 “Multilevel coarse graining of multiscale problems”, and Project C03 “Multiscale modelling and simulation for spatiotemporal master equation”).